Find partial z partial x and partial z partial y for


z ^ 4 - y ^ 2 + x ^ 2 = 6x ^ 2 * y ^ 3 * z ^ 7 .

1
Expert's answer
2022-01-11T19:13:36-0500

"z ^ 4 - y ^ 2 + x ^ 2 = 6x ^ 2 y ^ 3 z ^ 7"

Taking partial derivate w.r.t "x" both side for "z" ("y" will be treated as constant)

"4z^3\frac{\partial z}{\partial x}-0+2x=12xy^3z^7+42x^2y^3z^6 \frac{\partial z}{\partial x}\\
\Rightarrow \frac{\partial z}{\partial x}=\frac{12xy^3z^7}{4z^3-42x^2y^3z^6}"

Taking partial derivate w.r.t "y" both side for "z" ("x" will be treated as constant)

"4z^3\frac{\partial z}{\partial y}-2y+0=18x^2y^2 z^7+42x^2y^3z^6 \frac{\partial z}{\partial y}\\
\Rightarrow \frac{\partial z}{\partial y}=\frac{12xy^3z^7}{4z^3-18x^2y^2 z^7}"